OpenStudy (anonymous):

red:4 brown:10 yellow:7 green:9 blue:12 orange:13

OpenStudy (anonymous):

The primary colors are yellow, blue, and red. In simplest form, what fraction of the bag is made up of primary colors? Show your work and explain how you know.

OpenStudy (anonymous):

number of primary colors=7+12+4=23 total=55 fraction=23/55

OpenStudy (anonymous):

one more qeustion

OpenStudy (anonymous):

Do the primary colors make up more or less than half the total number of candies? Explain how the fraction you found in part a helps you know.

OpenStudy (anonymous):